# D-NPC: Dynamic Neural Point Clouds for Non-Rigid View Synthesis from Monocular Video

### v1

@@ Abstract
Dynamic reconstruction and spatiotemporal novel-view synthesis of non-rigidly deforming scenes recently gained increased attention. While existing work achieves impressive quality and performance on multi-view or teleporting camera setups, most methods fail to efficiently and faithfully recover motion and appearance from casual monocular captures. This paper contributes to the field by introducing a new method for dynamic novel view synthesis from monocular video, such as casual smartphone captures. Our approach represents the scene as a dynamic neural point cloud, an implicit time-conditioned point distribution that encodes local geometry and appearance in separate hash-encoded neural feature grids for static and dynamic regions. By sampling a discrete point cloud from our model, we can efficiently render high-quality novel views using a fast differentiable rasterizer and neural rendering network. Similar to recent work, we leverage advances in neural scene analysis by incorporating data-driven priors like monocular depth estimation and object segmentation to resolve motion and depth ambiguities originating from the monocular captures. In addition to guiding the optimization process, we show that these priors can be exploited to explicitly initialize our scene representation to drastically improve optimization speed and final image quality. As evidenced by our experimental evaluation, our dynamic point cloud model not only enables fast optimization and real-time frame rates for interactive applications, but also achieves competitive image quality on monocular benchmark sequences. Our project page is available at https://moritzkappel.github.io/projects/dnpc.
@@
figure We introduce Dynamic Neural Point Clouds for dynamic view synthesis from monocular videos. Our method enables fast optimization, real-time frame rates, and competitive image quality.
@@ 1 Introduction
Synthesizing novel views from a sparse set of input images is a fundamental challenge in computer vision. Using recent advances in 3D neural scene reconstruction, it becomes possible to create photorealistic renderings from arbitrary viewpoints, effectively adding interactive six degrees-of-freedom camera support. In the context of general, non-rigidly deforming dynamic scenes, the capability of interactively rendering novel views for every point in time enables numerous applications, including video stabilization [28], immersive playback for social media [2], or virtual reality [57]. However, performing faithful reconstruction from just a single monocular video (e.g., a casual smartphone recording) remains a computationally expensive and challenging problem.
@@
While the introduction of Neural Radiance Fields [31] and 3D Gaussian Splatting [19] brought about significant advances in static scene reconstruction, dynamic extensions still lag behind in terms of quality and robustness due to motion and depth ambiguities in the temporal domain. Those ambiguities are usually resolved using expensive large-scale multi view setups [24] or monocularized data (i.e., teleporting camera setups) [18], which maximize the ratio between camera and scene motion to improve temporal consistency. However, recent investigations [15] reveal that most sparsely regularized methods fail to faithfully recover consistent motion and geometry from casual smartphone captures due to the lack of effective multi-view signal. Multiple recent radiance field approaches address this problem by incorporating strong priors like monocular depth, optical flow, and semantic segmentation for the joint reconstruction of geometry, 4D scene flow fields [25, 26], and even camera parameters [27] from monocular video. While current monocular methods achieve high-quality spatiotemporally consistent view synthesis, they are often limited in the expressiveness of novel camera views, and by their high computational demand.
@@
To address the above-mentioned challenges, this paper introduces Dynamic Neural Point Clouds (D-NPC), a new system for monocular dynamic view synthesis based on implicit point distributions and fast hash-encoded feature grids. Given a monocular input video and the corresponding structure-from-motion calibration, our method optimizes a 4D representation of the depicted scene, which enables spatio-temporal novel-view synthesis, e.g., the infamous bullet-time effect. Following the concept of implicit neural point clouds [17], we track expected scene density in a sparse voxel grid, which is used to extract explicit pose-dependent point clouds. Combined with a fast multi-resolution hash-encoded feature grid [33], this scene representation is able to efficiently render high-fidelity novel views using a differentiable rasterizer and image-to-image translation network. Based on the concepts of space-time hash encoding [51], we lift the representation to the temporal domain by replacing the single feature grid with a 3D and 4D hash encoding for the static background and dynamic foreground regions, respectively. We further extend the point distribution field to robustly track the per-voxel ratio of static and dynamic scene content, effectively enabling 3D dynamics segmentation to significantly reduce the required compute during optimization and inference.
@@
To guide the severely under-constrained monocular reconstruction, we leverage per-frame data-driven priors, namely optical flow, monocular depth, and foreground segmentation masks, which we extract and preprocess with minimal overhead. We demonstrate that our temporal extension can take advantage of the rasterization-based forward rendering of implicit neural point clouds, which enables efficient foreground-background decomposition and image-space regularization. Furthermore, the scene-level sampling approach also provides a seamless way of integrating data-driven priors for model initialization, resulting in faster and better convergence during optimization. As shown in D-NPC: Dynamic Neural Point Clouds for Non-Rigid View Synthesis from Monocular Video, D-NPC achieves competitive image quality while significantly reducing optimization time, and allowing for interactive rendering in a graphical user interface.
@@ 2 Related Work
Rendering novel viewpoints of a recorded scene is a longstanding and well-studied problem. Early work on image-based rendering uses morphing to generate in-between views from image collections without explicit 3D reconstruction [9, 42]. Other approaches apply implicit or explicit 3D scene representations like light fields [16] or layered depth images [43] to generate high-quality novel views. Recently, Neural Radiance Fields (NeRFs) [31] and 3D Gaussian Splatting (3D-GS) [19] inspired a multitude of follow-up research, including extensions for dynamic scenes. We next briefly summarize neural representations for static scenes, followed by dynamic view synthesis.
@@ 2.1 Static View Synthesis
NeRFs [31] implicitly represent the scene geometry and appearance using a Multi-Layer Perceptron (MLP), yielding an unprecedented image quality for novel views at the cost of slow optimization and rendering due to the large number of network queries required. While some follow-ups propose qualitative improvements for, e.g., anti-aliasing and unbounded scenes [3, 4, 5], another branch of methods focuses on acceleration using explicit scene representations like sparse [13] or decomposed [8] voxel grids. Along the same lines, Müller et al. [33] propose an efficient grid-based scene representation implemented as a multi-resolution hash-grid, enabling high-quality reconstruction in only a few minutes. All the above volume rendering approaches apply backward rendering, where the scene is frequently sampled along the camera rays. In contrast, forward rendering approaches, explicitly model scene geometry using geometric primitives, allowing for fast rasterization-based rendering. The popular 3D-GS by Kerbl et al. [19] represents the scene as a set of 3D Gaussians for very fast rendering at a high image quality, inspiring many extensions [22, 60]. Analogously, point-based rendering approaches [1, 21, 39] use a fixed initial point cloud combined with a neural image translation network for hole filling. The recent INPC approach by Hahlbohm et al. [17] combines the benefits of forward and backward rendering, modelling the scene as a point probability and hash-encoded appearance field for high-quality rendering at interactive frame rates. Our D-NPC method leverages these benefits for fast dynamic view synthesis from monocular video. A more detailed discussion of neural rendering techniques can be found in the report by Tewari et al. [48].
@@ 2.2 Dynamic View Synthesis
Most NeRF-based extensions use a time-conditioned MLP [24, 25, 36] or additional deformation field [11, 34, 35, 37] to represent scene dynamics. Following acceleration models for static reconstruction, similar models based on factorized [2, 7, 12, 52] and hash-encoded [18, 45, 51] feature grids were introduced for dynamic view synthesis. D2NeRF [56] combined two models for foreground-background separation from a single video. Likewise, Wang et al. [51] decompose the scene in a fast 3D and 4D hash-grid representation. We adapt concepts from both approaches, integrating foreground/background handling in both our sampling and appearance modelling, enabling our method to individually render dynamic and static scene content while improving performance. Similar to radiance fields, recent approaches opt to extend 3D-GS for non-rigid scenes, tracking the 3D motion of Gaussians over time [29, 55] while retaining fast rendering. However, as recently revealed by Gao et al. [15], these methods require multi-view or monocularized [18] capturing setups to exploit observations from multiple views. On the other hand, multiple methods focus on dynamic view synthesis from a single monocular input video, integrating additional priors like dense scene flow estimation to resolve motion and depth ambiguities [25, 26, 27]. RoDynRF [27] further takes the challenging scenario to the extreme, additionally optimizing camera poses from the unposed input video. However, these MLP-based monocular approaches rely on backward rendering, thus being limited in optimization speed and final rendering frame rate. This paper presents a point-based forward rendering approach for monocular view synthesis, which leverages additional priors to initialize and guide the reconstruction, enabling high-quality rendering at real-time frame rates. More details on the state of the art in dynamic reconstruction can be found in the survey by Yunus et al. [61].
@@ 3 Method
We introduce Dynamic Neural Point Clouds (D-NPC), a method for dynamic novel-view synthesis from monocular video. Given a single monocular video consisting of $N$ images $I=(i_{1},...,i_{N})$ and their normalized timestamps $T=(t_{1},...,t_{N})\in[0,1]^{N}$, we reconstruct a dynamic implicit point cloud representation [17] of the depicted dynamic scene. Based on its static counterpart, the core of our D-NPC model consists of a spatiotemporal point probability ($\mathcal{P}$) and appearance feature descriptor ($\mathcal{A}$) field, which can efficiently be rendered using a differentiable rasterizer and neural rendering network. A full overview of our method is given in Fig. 1. In the following, we describe the individual components and optimization process in detail.
@@ 3.1 Data Preprocessing
The reconstruction of non-rigidly deforming scenes from a single moving camera is a highly ill-posed problem. To regularize the reconstruction from monocular video, previous approaches [15, 25, 27] apply estimates from several image processing applications, replacing the multi-view consistency signal with learned, data-driven scene analysis. Following this idea, our method starts with data preprocessing to extract high-quality priors, which we use to initialize and guide our model in the later stages. Despite the recent emergence of consistent yet expensive scene calibration [44], video depth estimation [30] and motion tracking [53] estimators, one of our method’s main objectives is speed and simplicity. Thus, we focus on real-time out-of-the-box solutions to maintain a fast total reconstruction time (including prior computation).
First, we start by applying sparse structure-from-motion (SfM) using COLMAP [40, 41] to estimate camera intrinsics and per-frame poses (extrinsics) $E=(e_{1},...,e_{N})$. As a side product, we extract the sparse, globally-aligned SfM point cloud. We further use RAFT [47] to estimate dense optical flow in the forward and backward directions, and estimate per frame (dis-)occlusion masks based on forward-backward consistency. To reduce the depth ambiguity of dynamic scene content, we employ the fast monocular depth estimator DepthAnything [58], which provides relative disparity maps for every video frame. Instead of using a relative depth loss [25], we align the predicted disparity with the global SfM point cloud by estimating a scale and shift via RANSAC-based least squares and temporal smoothing. This way, we obtain globally-aligned depth which can explicitly be integrated into our model. Lastly, we obtain binary segmentation masks $\omega$ of moving foreground object(s) to mask out dynamic scene content in our reconstruction and COLMAP SfM feature extraction. Similar to monocular depth, automatic motion segmentation is highly ill-posed, even in the presence of semantic segmentation masks. Thus, we use the semi-automatic video object segmentation system Cutie [10], which is guided by minimal user input in the form of manual keypoint selection. In line with our demands on simplicity, Cutie generates masks of sufficient quality in less than a minute, requiring as few as two mouse clicks, which in the future could be integrated into the data recording setup on mobile devices.
@@ 3.2 Dynamic Neural Point Cloud Representation
Our method represents the 4D scene as a Dynamic Neural Point Cloud, the first temporal extension of implicit neural point clouds [17]. The primary idea of implicit point clouds is to replace the persistent point cloud $(p,f)$, consisting of a set of 3D positions $p$ and corresponding color or feature descriptor vectors $f$, with a sampling distribution $\mathcal{P}$ and a neural feature grid $\mathcal{A}$. By repeatedly sampling and rendering an explicit point cloud $((p,\mathcal{A}(p)),p\sim\mathcal{P})$, it is possible to iteratively update and refine the spatial distribution, while maintaining all benefits of point-based rendering. However, applying this concept to dynamic scenes is not straightforward, as the direct integration of a fourth dimension is memory-intensive and does not naturally exploit any temporal coherency.

### v2

@@ Abstract
Dynamic reconstruction and spatiotemporal novel-view synthesis of non-rigidly deforming scenes recently gained increased attention. While existing work achieves impressive quality and performance on multi-view or teleporting camera setups, most methods fail to efficiently and faithfully recover motion and appearance from casual monocular captures. This paper contributes to the field by introducing a new method for dynamic novel view synthesis from monocular video, such as casual smartphone captures. Our approach represents the scene as a dynamic neural point cloud, an implicit time-conditioned point distribution that encodes local geometry and appearance in separate hash-encoded neural feature grids for static and dynamic regions. By sampling a discrete point cloud from our model, we can efficiently render high-quality novel views using a fast differentiable rasterizer and neural rendering network. Similar to recent work, we leverage advances in neural scene analysis by incorporating data-driven priors like monocular depth estimation and object segmentation to resolve motion and depth ambiguities originating from the monocular captures. In addition to guiding the optimization process, we show that these priors can be exploited to explicitly initialize our scene representation to drastically improve optimization speed and final image quality. As evidenced by our experimental evaluation, our dynamic point cloud model not only enables fast optimization and real-time frame rates for interactive applications, but also achieves competitive image quality on monocular benchmark sequences.
Our code and data are available online https://moritzkappel.github.io/projects/dnpc/.
{CCSXML}
<ccs2012> <concept> <concept_id>10010147.10010371.10010382.10010385</concept_id> <concept_desc>Computing methodologies Image-based rendering</concept_desc> <concept_significance>100</concept_significance> </concept> <concept> <concept_id>10010147.10010371.10010396.10010400</concept_id> <concept_desc>Computing methodologies Point-based models</concept_desc> <concept_significance>500</concept_significance> </concept> <concept> <concept_id>10010147.10010178.10010224.10010245.10010254</concept_id> <concept_desc>Computing methodologies Reconstruction</concept_desc> <concept_significance>100</concept_significance> </concept> <concept> <concept_id>10010147.10010371.10010372.10010373</concept_id> <concept_desc>Computing methodologies Rasterization</concept_desc> <concept_significance>300</concept_significance> </concept> </ccs2012>
@@
[…]
@@ 1 Introduction
Synthesizing novel views from a sparse set of input images is a fundamental challenge in computer vision. Using recent advances in 3D neural scene reconstruction, it becomes possible to create photorealistic renderings from arbitrary viewpoints, effectively adding interactive six degrees-of-freedom camera support. In the context of general, non-rigidly deforming dynamic scenes, the capability of interactively rendering novel views for every point in time enables numerous applications, including video stabilization [LLY21], immersive playback for social media [AHR23], or virtual reality [XAL23]. However, performing faithful reconstruction from just a single monocular video (e.g., a casual smartphone recording) remains a computationally expensive and challenging problem.
@@
While the introduction of Neural Radiance Fields [MST21] and 3D Gaussian Splatting [KKLD23] brought about significant advances in static scene reconstruction, dynamic extensions still lag behind in terms of quality and robustness due to motion and depth ambiguities in the temporal domain. Those ambiguities are usually resolved using expensive large-scale multi view setups [LSZ22] or monocularized data (i.e., teleporting camera setups) [KGC25], which maximize the ratio between camera and scene motion to improve temporal consistency. However, recent investigations [GLT22] reveal that most sparsely regularized methods fail to faithfully recover consistent motion and geometry from casual smartphone captures due to the lack of effective multi-view signal. Multiple recent radiance field approaches address this problem by incorporating strong priors like depth, optical flow, and semantic segmentation for the joint reconstruction of geometry, 4D scene flow fields [LNSW21, LWC23], and even camera parameters [LGM23] from monocular video. While current monocular methods achieve high-quality spatio-temporally consistent view synthesis, they are often limited in the expressiveness of novel camera views, and by their high computational demand.
@@
To address the above-mentioned challenges, this paper introduces Dynamic Neural Point Clouds (D-NPC), a new system for monocular dynamic view synthesis based on implicit point distributions and fast hash-encoded feature grids. Given a monocular input video and the corresponding structure-from-motion calibration, our method optimizes a 4D representation of the depicted scene, which enables spatio-temporal novel-view synthesis, e.g., the infamous bullet-time effect. Following the concept of implicit neural point clouds [HFK24], we track expected scene density in a sparse voxel grid, which is used to extract explicit pose-dependent point clouds. Combined with a fast multi-resolution hash-encoded feature grid [MESK22], this scene representation is able to efficiently render high-fidelity novel views using a differentiable rasterizer and image-to-image translation network. Based on the concepts of space-time hash encoding [WCW23], we lift the representation to the temporal domain by replacing the single feature grid with a 3D and 4D hash encoding for the static background and dynamic foreground regions, respectively. We further extend the point distribution field to robustly track the per-voxel ratio of static and dynamic scene content, effectively enabling 3D dynamics segmentation to significantly reduce the required compute during optimization and inference.
@@
To guide the severely under-constrained monocular reconstruction, we leverage per-frame data-driven priors, namely monocular depth and foreground segmentation masks, which we extract and preprocess with minimal overhead. We demonstrate that our temporal extension can take advantage of the rasterization-based forward rendering of implicit neural point clouds, which enables efficient foreground-background decomposition and image-space regularization. Furthermore, the scene-level sampling approach also provides a seamless way of integrating data-driven priors for model initialization, resulting in faster and better convergence during optimization. As shown in D-NPC: Dynamic Neural Point Clouds for Non-Rigid View Synthesis from Monocular Video, D-NPC achieves competitive image quality while significantly reducing optimization time, and allowing for interactive rendering in a graphical user interface.
@@ 2 Related Work
Rendering novel viewpoints of a recorded scene is a longstanding and well-studied problem. Early work on image-based rendering uses morphing to generate in-between views from image collections without explicit 3D reconstruction [CW93, SD96]. Other approaches apply implicit or explicit 3D scene representations like light fields [GGSC96], per-view meshes [HPP18], geometric scaffolds [RK21], or layered depth images [SGHS98] to generate high-quality novel views. Recently, Neural Radiance Fields (NeRFs) [MST21] and 3D Gaussian Splatting (3D-GS) [KKLD23] inspired a multitude of follow-up research, including extensions for dynamic scenes. We next briefly summarize neural representations for static scenes, followed by dynamic view synthesis.
@@ 2.1 Static View Synthesis
NeRFs [MST21] implicitly represent the scene geometry and appearance using a Multi-Layer Perceptron (MLP), yielding an unprecedented image quality for novel views at the cost of slow optimization and rendering due to the large number of network queries required. While some follow-ups propose qualitative improvements for, e.g., anti-aliasing and unbounded scenes [BMT21, BMV22, BMV23], another branch of methods focuses on acceleration using explicit scene representations like sparse [FKYT22] or decomposed [CXG22] voxel grids. Along the same lines, Müller et al. [MESK22] propose an efficient grid-based scene representation implemented as a multi-resolution hash-grid, enabling high-quality reconstruction in only a few minutes. All the above volume rendering approaches apply backward rendering, where the scene is frequently sampled along the camera rays. In contrast, forward rendering approaches, explicitly model scene geometry using geometric primitives, allowing for fast rasterization-based rendering. The popular 3D-GS by Kerbl et al. [KKLD23] represents the scene as a set of 3D Gaussians for very fast rendering at high image quality, inspiring many extensions [LRS24, YCH24]. Analogously, point-based rendering approaches [ASK20, KPLD21, RFS22] use a fixed initial point cloud combined with a neural image translation network for hole filling. The recent INPC approach by Hahlbohm et al. [HFK24] combines the benefits of forward and backward rendering, modelling the scene as a point probability and hash-encoded appearance field for high-quality rendering at interactive frame rates. Our D-NPC method leverages these benefits for fast dynamic view synthesis from monocular video. A more detailed discussion of neural rendering techniques can be found in the report by Tewari et al. [TTM22].
@@ 2.2 Dynamic View Synthesis
Most NeRF-based extensions use a time-conditioned MLP [LSZ22, LNSW21, PSJ23] or additional deformation field [FYW22, PSB21, PSH21, PCPMMN21] to represent scene dynamics. Following acceleration models for static reconstruction, similar models based on factorized [AHR23, CJ23, FKMW23, WTL23] and hash-encoded [KGC25, SCL23, WCW23] feature grids were introduced for dynamic view synthesis. D2NeRF [WZT22] combined two models for foreground-background separation from a single video. Likewise, Wang et al. [WCW23] decompose the scene in a fast 3D and 4D hash-grid representation. We adapt concepts from both approaches, integrating foreground/background handling in both our sampling and appearance modelling, enabling our method to individually render dynamic and static scene content while improving performance. Similar to radiance fields, recent approaches opt to extend 3D-GS for non-rigid scenes, tracking the 3D motion of Gaussians over time [LKLR24, WYF24] while retaining fast rendering. However, as recently revealed by Gao et al. [GLT22], these methods require multi-view or monocularized [KGC25] capturing setups to exploit observations from multiple views. On the other hand, multiple methods focus on dynamic view synthesis from a single monocular input video, integrating additional priors like dense scene flow estimation to resolve motion and depth ambiguities [LNSW21, LWC23, LGM23]. RoDynRF [LGM23] further takes the challenging scenario to the extreme, additionally optimizing camera poses from the unposed input video. However, these MLP-based monocular approaches rely on backward rendering, thus being limited in optimization speed and final rendering frame rate. To circumvent slow optimization, the PGDVS approach by Zhao et al. [ZCM24] proposes a generalized approach for monocular video that bypasses per scene optimization. Still, the slow inference speed and varying generalization quality largely rule out application for real-time downstream tasks.
To jointly enable fast high quality per-scene optimization and subsequent real-time novel-view synthesis, this paper presents a point-based forward rendering approach for monocular view synthesis, which leverages additional priors to initialize and guide the reconstruction. More details on the state of the art in dynamic reconstruction can be found in the survey by Yunus et al. [YLN24].
2.3 Concurrent Work
Concurrent to our work, other approaches propose fast forward rendering representations for dynamic reconstruction and novel-view synthesis from monocular video. Please note that the covered concurrent methods are not yet published in any proceedings (at the time of submission), or constitute non-peer reviewed work on arXiv.
Similar to our approach, Lee et al. [LZBM25] separately model the foreground and background, but instead use a plane based proxy geometry for static regions, and points for the dynamic foreground, enabling fast forward rendering at a decent image quality. Following up on the adaption of 3DGS for dynamic multi-view recordings, Stearns et al. [SHU24] propose a divide-and-conquer approach for motion and appearance reconstruction from single monocular video using isotropic Gaussians. Like most Gaussian-based methods, this approach inherits impressive online framerates and decent image quality, but adds overhead in terms of optimization time. The Shape of Motion approach by Wang et al. [WYG24] additionally uses 2D motion track inputs to recover 3D scene motion represented via decomposed motion bases, which is applied to a set of canonical 3D Gaussians for novel view synthesis. Lei et al. [LWH24] also propose to disentangle the deformations of the scene from its geometry. The former is represented as a sparse graph of motion trajectories that they lift from 2D priors and optimize exploiting semantic features and the assumption of the deformations being as-rigid-as-possible. They use Gaussian Splatting to render the scene resulting from the global fusion of the 3DGs per timestep and deformed accordingly to the motion graph. While our method also involves rendering explicit geometry for improved performance, using an implicit point cloud representation instead of explicit Planes, Points, or Gaussians entails unique characteristics regarding performance, quality and robustness.
@@ 3 Method
We introduce Dynamic Neural Point Clouds (D-NPC), a method for dynamic novel-view synthesis from monocular video. Given a single monocular video consisting of $N$ images $I=(i_{1},...,i_{N})$ and their normalized timestamps $T=(t_{1},...,t_{N})\in[0,1]^{N}$, we reconstruct a dynamic implicit point cloud representation [HFK24] of the depicted dynamic scene. Based on its static counterpart, the core of our D-NPC model consists of a spatiotemporal point probability ($\mathcal{P}$) and appearance feature descriptor ($\mathcal{A}$) field, which can efficiently be rendered using a differentiable rasterizer and neural rendering network. A full overview of our method is given in Fig. 1. In the following, we describe the individual components and optimization process in detail.
@@ 3.1 Data Preprocessing
The reconstruction of non-rigidly deforming scenes from a single moving camera is a highly ill-posed problem. To regularize the reconstruction from monocular video, previous approaches [GLT22, LNSW21, LGM23] apply estimates from several image processing applications, replacing the multi-view consistency signal with learned, data-driven scene analysis. Following this idea, we first extract additional priors, which we use to initialize and guide our model in the later stages. Despite the recent emergence of consistent yet expensive scene calibration [SCTS24], video depth estimation [LHS20] and motion tracking [WCC23] estimators, one of our method’s main objectives is speed and simplicity. Thus, we focus on fast out-of-the-box solutions to maintain a swift total reconstruction time (including prior computation).
First, we start by applying sparse structure-from-motion (SfM) using COLMAP [SF16, SZFP16] or VGGSfM [WKRN24] to estimate camera intrinsics and per-frame poses (extrinsics) $E=(e_{1},...,e_{N})$. As a side product, we extract the sparse, globally-aligned SfM point cloud. To reduce the depth ambiguity of dynamic scene content, we employ the fast monocular depth estimator DepthAnything [YKH24], which provides relative disparity maps for every video frame. Instead of using a relative depth loss [LNSW21], we align the predicted disparity with the global SfM point cloud by estimating a scale and shift via RANSAC-based least squares and temporal smoothing. This way, we obtain globally-aligned depth which can explicitly be integrated into our model. Alternatively, our method also accepts (sparse) metric LiDAR depth and camera poses obtained, e.g., via IMUs, which completely lifts the dependency on SfM calibration and point cloud priors. Lastly, we obtain binary segmentation masks $\omega$ of moving foreground object(s) to mask out dynamic scene content in our reconstruction and COLMAP SfM feature extraction. Similar to monocular depth, automatic motion segmentation is highly ill-posed, even in the presence of semantic segmentation masks. Thus, we use the semi-automatic video object segmentation system Cutie [COP24], which is guided by minimal user input. In line with our demands on simplicity, Cutie generates masks of sufficient quality in less than a minute, requiring as few as two mouse clicks, which in the future could be integrated into the data recording setup on mobile devices.
@@ 3.2 Dynamic Neural Point Cloud Representation
Our method represents the 4D scene as a Dynamic Neural Point Cloud, the first temporal extension of implicit neural point clouds [HFK24]. In their paper, Hahlbohm et al. propose a methodology that diverges from the conventional approach of storing both the spatial and photometric properties of the scene within a unified data structure. Instead, they advocate for a division of these properties and an implicit optimization of both. The geometric structure is depicted as an octree-based point probability field ($\mathcal{P}$), which undergoes progressive subdivision to ensure a uniform probability distribution across all leaf nodes. Conversely, the appearance features are embedded in an implicit coordinate-based multi-resolution hash grid ($\mathcal{A}$). During rendering, $\mathcal{P}$ is employed as an estimator for point positions, utilizing either random positions in each leaf or fixed sampling patterns. Meanwhile, per-point appearance features are queried from $\mathcal{A}$. The authors then employ fast bilinear splatting to render the resulting explicit point cloud, where gradients are backpropagated through their pipeline to the implicit representation [HFK24].
Thus, the primary idea of implicit point clouds is to replace the persistent point cloud $(p,f)$, consisting of a set of 3D positions $p$ and corresponding color or feature descriptor vectors $f$, with a sampling distribution $\mathcal{P}$ and a neural feature grid $\mathcal{A}$. By repeatedly sampling and rendering an explicit point cloud $((p,\mathcal{A}(p)),p\sim\mathcal{P})$, it is possible to iteratively update and refine the spatial distribution, while maintaining all benefits of point-based rendering. However, applying this concept to dynamic scenes is not straightforward, as the direct integration of a fourth dimension is memory-intensive and does not naturally exploit any temporal coherency.